### v1

# Mitigating the Accuracy-Robustness Trade-off via Multi-Teacher Adversarial Distillation

Abstract
Adversarial training is a practical approach for improving the robustness of deep neural networks against adversarial attacks. Although bringing reliable robustness, the performance toward clean examples is negatively affected after adversarial training, which means a trade-off exists between accuracy and robustness. Recently, some studies have tried to use knowledge distillation methods in adversarial training, achieving competitive performance in improving the robustness but the accuracy for clean samples is still limited. In this paper, to mitigate the accuracy-robustness trade-off, we introduce the Multi-Teacher Adversarial Robustness Distillation (MTARD) to guide the model’s adversarial training process by applying a strong clean teacher and a strong robust teacher to handle the clean examples and adversarial examples, respectively. During the optimization process, to ensure that different teachers show similar knowledge scales, we design the Entropy-Based Balance algorithm to adjust the teacher’s temperature and keep the teachers’ information entropy consistent. Besides, to ensure that the student has a relatively consistent learning speed from multiple teachers, we propose the Normalization Loss Balance algorithm to adjust the learning weights of different types of knowledge. A series of experiments conducted on public datasets demonstrate that MTARD outperforms the state-of-the-art adversarial training and distillation methods against various adversarial attacks.
Index Terms:
DNNs, Adversarial Training, Knowledge Distillation, Adversarial Robustness, Accuracy-Robustness Trade-off.
1 Introduction
Deep Neural Networks (DNNs) have become powerful tools for solving complex real-world learning problems, such as image classification [13], face recognition [34], and natural language processing [29]. However, Szegedy et al. [32] demonstrate that DNNs are vulnerable to adversarial attacks with imperceptible adversarial perturbations on input, which causes wrong predictions of DNNs. To defend against adversarial attacks, adversarial training is proposed and has shown its effectiveness in obtaining adversarial robust DNNs [35, 6, 25]. While improving the robustness of DNNs, a negative impact exists on the model’s accuracy on clean samples. Some methods [45, 26] are proposed to alleviate the trade-off between accuracy and robustness. However, this phenomenon still exists and needs to be further explored.
Recently, to further enhance the robustness of small DNNs, Knowledge Distillation [15] is applied as a powerful tool in adversarial training. Knowledge Distillation can transfer knowledge from strong robust models to the student model, which utilizes the predictions from the teacher models as label information to guide the student model [9, 48, 49]. These methods are called adversarial robustness distillation. Compared with one-hot ground truth labels, the teacher’s predicted labels can not only retain the correctness of the target label but also reflect richer knowledge information of the non-target label. Several studies attempt to explain why knowledge distillation is effective through various views, e.g., Re-weighting [8], Privileged Information [24], and Label-smoothing [43]. Recently, Li et al. [20] argue that correctness, smooth regularization, and class discriminability in teacher’s predicted labels can make up the knowledge of the teacher’s predicted distribution, and that appropriate category discriminability plays an important role in improving the effectiveness of knowledge distillation. While achieving impressive robust performance, we are curious if adversarial robustness distillation could address the trade-off between accuracy and robustness.
Generally speaking, if we want to improve both accuracy and robustness through knowledge distillation, the provided teacher model is supposed to perform well in both aspects. That is to say, the teacher model should equally provide correct guidance and rich knowledge information for clean and adversarial examples. However, due to the existing trade-off, a single teacher model is difficult to achieve the above goals. Therefore, a reasonable approach is to decouple adversarial knowledge distillation, i.e., using two teachers that are good at accuracy and robustness respectively (they can be called the clean teacher and the robust teacher) to guide the student in different knowledge types. Under this divide-and-conquer strategy, the student model guided by multiple teacher models can enhance its performance in both aspects theoretically.
However, the actual implementation of this idea is challenging. Owing to the intrinsic trade-off between accuracy and robustness, the clean teacher and robust teacher push the student model towards opposite directions. This is obviously different from the current multi-teacher knowledge distillation methods [31, 42, 23], where their optimization goal of different teachers is consistent, i.e., all teachers push the student towards the good clean accuracy. The trade-off in our task leads to a difficult optimization mainly in two aspects: (1) The two teacher models should have similar knowledge scales themselves. When one teacher has too much knowledge scale, the student will be forced to obtain more knowledge from this teacher and relatively less knowledge from another, which directly results in an imbalance between accuracy and robustness. However, finding a clean teacher and a robust teacher with similar knowledge scales is not easy. Due to the different network architecture and training methods, a natural gap exists in knowledge scales between different teachers. Meanwhile, how to accurately measure teachers’ knowledge scales is also an unsolved problem. (2) The student’s learning ability from different types of knowledge should be balanced. More specifically, because two hugely different optimization directions exist in adversarial robustness distillation, the student may over-fit an ability that is easier to learn but ignore another due to different learning difficulties. Only the cooperation of the student and teachers can help the student obtain both strong accuracy and robustness.
Based on the above discussions, in this paper, we propose the Multi-Teacher Adversarial Robustness Distillation (MTARD) to enhance both accuracy and robustness at the same time. During adversarial training, a clean teacher is to instruct the student for handling clean examples and a strong robust teacher is to instruct the student for handling adversarial examples. To meet the two challenges above, the novel Entropy-Based Balance and Normalization Loss Balance strategies are presented.
For the first challenge, to ensure that the two teacher models show similar knowledge scales during the optimization, we propose an Entropy-Based Balance algorithm to balance the information entropy among different teachers’ predicted distribution. Specifically, based on information theory [30], relative entropy can measure the incremental information from an initialized network’s predicted distribution to the well-trained teachers’ predicted distribution, so we utilize the relative entropy to measure the teachers’ knowledge scales. Furthermore, we theoretically prove that the difference in knowledge scales can be transformed into the difference in teachers’ information entropy. Thus Entropy-Based Balance algorithm can balance the knowledge scales by adjusting each teacher’s temperature until the information entropy is appropriate and consistent.
For the second challenge, to maintain the relative equality of the student’s learning speed from multiple teachers, we propose a Normalization Loss Balance algorithm to control loss weight and balance the influence between the clean teacher and the robust teacher. Specifically, inspired by [4], we design a relative loss to measure the knowledge proportion that the student learns from different teachers. The different loss weights for the update of student parameters are dynamically adjusted to keep the relative loss consistent during the training process. Compared with the initial state, the student model will be forced to learn relatively equal knowledge scales from all the teacher models.
The main contributions of this work are three-fold:
  • We propose a novel framework: Multi-Teacher Adversarial Robustness Distillation (MTARD). MTARD decouples the adversarial knowledge distillation to obtain accuracy and robustness respectively. We apply a clean teacher and a robust teacher to bring both clean and robust knowledge to the student.
  • We propose the Entropy-Based Balance algorithm and Normalization Loss Balance algorithm to balance multiple teachers. The Entropy-Based Balance algorithm is applied to control the teachers’ knowledge scales by dynamically adjusting temperatures. The Normalization Loss Balance algorithm is utilized to ensure that the student has relatively equal learning speeds for different teachers’ knowledge.
  • We empirically verify the effectiveness of MTARD in improving performance. The Weighted Robust Accuracy (a measure to evaluate both clean and robust accuracy) of our MTARD-trained models improves significantly against a variety of attacks compared to state-of-the-art adversarial training and knowledge distillation methods. Especially for query-based black-box attacks, MTARD can enhance the Weighted Robust Accuracy by 5.56%, 4.78%, and 3.31% for MobileNet-v2 on CIFAR-10, CIFAR-100, and Tiny-ImageNet, respectively.
This journal paper is an extended version of our ECCV paper [47]. Compared with the conference version, we have made significant improvements and extensions in this version in the following aspects: (1) At the idea level, we discuss the trade-off problem in a more comprehensive view from both teachers and the student, and then point out that the difference in the two teachers’ knowledge scales will affect the trade-off between accuracy and robustness (Section 1). (2) At the method level, a multiple teacher’s knowledge adjustment mechanism named the Entropy-Based Balance algorithm is designed to implement the new idea (Section 3.2). (3) At the experimental level, we additionally train models on the Tiny-ImageNet dataset and compare our proposed method with more SOTA methods against more advanced attacks. In addition, we give more ablation studies to comprehensively test our method (Section 4). (4) At the writing level, the introduction, related works, methods, and experiments have been rewritten and refined.
The rest of the paper is organized as follows: We introduce the related work in Section 2. Section 3 proposes our MTARD to enhance both accuracy and robustness. The experiments are conducted in Section 4, and the conclusion is given in Section 5.
2 Related Work
2.1 Adversarial Attacks
Since Szegedy et al. [32] propose that adversarial examples can mislead the deep neural network, lots of effective adversarial attack methods, such as the Fast Gradient Sign Method (FGSM) [10], Projected Gradient Descent Attack (PGD) [25], and Carlini and Wagner Attack (CW) [3] are proposed. Attack methods can be divided into white-box attacks and black-box attacks. White-box attacks know all the parameter information of the target model when generating adversarial examples, and black-box attacks know little or no information. In general, black-box attacks simulate the model gradient by repeatedly querying the target model (query-based attack) [1, 39, 36, 37, 38] or searching for an alternative model similar to the target model (transfer-based attack) [7, 16]. Recently, a strong ensemble adversarial attack method named AutoAttack (AA) [6] is proposed, which consists of four attack methods, including Auto-PGD (APGD), Difference of Logits Ratio (DLR) attack, FAB-Attack [5], and the black-box Square Attack [1].
2.2 Adversarial Training
Adversarial Training [25, 45] is seen as an effective way to defend against adversarial attacks. Madry et al. [25] formulate Adversarial Training as a min-max optimization problem as follows:
$\min\limits_{\theta}E_{(x,y)\sim\mathcal{D}}[\max\limits_{\delta\in\Omega}\mathcal{L}(f(x+\delta;\theta),y)],$ (1)
where $f$ represents a deep neural network with weight $\theta$, $D$ represents a distribution of the clean example $x$ and the ground truth label $y$. $\mathcal{L}$ represents the loss function. $\delta$ represents the adversarial perturbation, and $\Omega$ represents a bound, which can be defined as $\Omega=\left\{\delta:||\delta||\leq\epsilon\right\}$ with the maximum perturbation scale $\epsilon$.
Much work has been proposed to further improve the robustness of adversarial training. Zhang et al. [45] attempt to reduce the gap between accuracy and robustness by minimizing a Kullback–Leibler (KL) divergence loss (TRADES). Wang et al. [35] further improve performance through Misclassification-Aware Adversarial Training (MART). Pang et al. [26] propose to employ local equivariance to describe the ideal behavior of a robust model, which facilitates the reconciliation between accuracy and robustness (SCORE).
Previous work [45, 26] uses two indexes: clean and robust accuracy, as the metric to evaluate the model’s performance. Gurel et al. [12] propose a metric named Weighted Robust Accuracy to comprehensively measure the model’s performance, which is used in our experiments.
2.3 Knowledge Distillation
Knowledge distillation can transfer the performance of other models to the target model [15]. Extensive research has been widely studied in recent years [18, 42, 20]. Knowledge distillation can briefly be formulated as the following optimization:
$\mathop{\arg\min}_{\theta_{S}}(1-\alpha)CE(S(x),y)+\alpha\tau^{2}KL(S(x;\tau),T(x;\tau)),$ (2)
where $KL$ is Kullback–Leibler divergence loss, $CE$ represents the cross-entropy loss. $\tau$ is a temperature constant combined with softmax operation, and the prediction $p_{k}$ with temperature $\tau$ can be formulated as follows:
$p_{k}=\frac{exp(z_{k}(x)/\tau)}{\sum_{j=1}^{C}exp(z_{j}(x)/\tau)},$ (3)
where $z_{k}(x)$ denotes the $k$-th dimension output logits of model before softmax layer, and $p_{k}$ denotes the predicted possibility of $k$-th class.
The temperature $\tau$ is artificially adjusted as a hyper-parameter in previous work. Recently, [21, 22] propose to adjust $\tau$ automatically based on the designed learnable sub-network. These methods need to train the sub-network according to the feedback of the student model, which adds additional training overhead. Our method automatically adjusts the temperature only based on multiple teachers’ knowledge scales and has almost no computational cost.
Some previous studies exist on multi-teacher knowledge distillation from different views: including response-based knowledge [18, 42], feature-based knowledge [23, 46], and relation-based knowledge [41, 40]. Previous methods often use multiple teachers for knowledge distillation for the same examples, and those teachers tend to optimize the student in similar directions. Different from previous methods, our multi-teacher adversarial robustness distillation targets two different kinds of examples. Since adversarial examples are generated to mislead the model, there is a huge gap between the optimization direction of clean examples and adversarial examples, which exist great difficulties compared with previous research.
2.4 Adversarial Robustness Distillation
To defend against the adversarial attack, the defense distillation [27] is proposed to use the knowledge distillation to defend against adversarial attack. However, this method is considered as “Obfuscated Gradients” [2] and can be broken by CW attack [3]. Then a series of methods [49, 9, 48] apply knowledge distillation to further enhance the robustness based on adversarial training [25]. ARD [9] proposes that using a strong robust model as the teacher model in adversarial training can achieve pretty robustness. IAD [48] performs adversarial knowledge distillation by compositing with unreliable teacher guidance and student introspection. RSLAD [49] uses the soft labels generated by the robust teacher model to produce adversarial examples, and further utilizes the robust labels to guide the training process of both clean examples and adversarial examples, which can effectively improve the robustness of the student.
The above methods focus on improving the robustness but ignore the performance of identifying clean examples. In our method, we consider both accuracy and robustness through clean and robust teachers, which comprehensively improves the performance of the student model.
3 Methodology
In this section, we propose MTARD to enhance both accuracy and robustness by applying the clean teacher and the robust teacher to guide the student; From the teachers’ teaching view, the Entropy-Based Balance algorithm is proposed to balance the teacher’s knowledge scale. From the student’s learning view, the Normalization Loss Balance algorithm is designed to balance the student’s learning speeds from different teachers.
Figure 1: The framework of our Multi-Teacher Adversarial Robustness Distillation (MTARD). In the training process, we first generate adversarial examples from the student. Then, with the guidance of the clean teacher and the robust teacher, the student is trained on clean examples and adversarial examples, respectively. Here, we apply the Entropy-Based Balance algorithm to adjust the teacher’s knowledge scales; In addition, we use the Normalization Loss Balance algorithm to balance the student’s knowledge learning speed from different teachers.
3.1 Multi-Teacher Adversarial Robustness Distillation
The training process in MTARD can be considered as a min-max optimization based on adversarial training. The whole framework can be viewed in Fig. 1. In the inner maximization, the adversarial examples are generated by the student based on the clean examples. In the outer minimization, the clean teacher and the robust teacher generate predicted labels to guide the student training process toward the clean examples and the adversarial examples, respectively. The optimization of the basic MTARD is defined as follows:
$\begin{split}\mathop{\arg\min}_{\theta_{S}}(1-\alpha)&KL(S(x_{nat};\tau_{s}),T_{nat}(x_{nat};\tau_{nat}))\\ &+\alpha KL(S(x_{adv};\tau_{s}),T_{adv}(x_{adv};\tau_{adv})),\end{split}$ (4)
$x_{adv}=\mathop{\arg\max}_{\delta\in\Omega}CE(S(x_{nat}+\delta),y),$ (5)
where $x_{nat}$ and $x_{adv}$ are clean examples and adversarial examples, $S(x;\tau_{s})$ represents student network $S$ with temperature $\tau_{s}$. $T_{nat}(x;\tau_{nat})$ and $T_{adv}(x;\tau_{adv})$ represent the clean and robust teacher with temperature $\tau_{nat}$ and $\tau_{adv}$, respectively. $\alpha$ is a constant in the basic proposal. $CE$ denotes the Cross-Entropy loss in the maximization. Here, $\tau_{s}$, $\tau_{nat}$, and $\tau_{adv}$ are usually pre-defined in previous works.
The goal of MTARD is to learn a student with strong accuracy as the clean teacher and strong robustness as the robust teacher. In the actual operation process, however, two teachers may display different knowledge scales; Meanwhile, the student may have different learning speeds toward a different type of teacher’s knowledge. The above two points may lead to an unbalanced performance in accuracy and robustness. So how to handle multi-teacher adversarial robustness distillation becomes a problem to be solved in the following two subsections.
3.2 Entropy-Based Balance in MTARD
We can theoretically obtain a student with both accuracy and robustness by introducing two teachers. However, due to the different model architectures and training methods, the teachers have different knowledge scales. Eliminating the trade-off between accuracy and robustness will become unrealistic if the student learns from the two teachers with no equal knowledge scales. The student will enjoy simultaneously acquiring equal knowledge scales from two teachers who appear to be equally knowledgeable. Based on the above analysis, we propose the Entropy-Based Balance algorithm in MTARD.
Then how to quantify the teacher’s knowledge scale needs to be solved. Intuitively speaking, a well-trained network model is trained from a network model with randomly initialized weights. The model with randomly initialized weights can be considered as a general model $I$ without any knowledge; After training, the model can be considered to possess the knowledge brought by the different optimization processes. Based on information theory, relative entropy (also known as Kullback–Leibler divergence [30]) represents the information cost required from one distribution to another distribution. Therefore, we apply the relative entropy to denote clean teacher’s knowledge scales $K_{nat}^{T_{nat}}$ and robust teacher’s knowledge scales $K_{adv}^{T_{adv}}$ as follows:
$K_{nat}^{T_{nat}}=KL(I(x_{nat}),T_{nat}(x_{nat})),$ (6)
$K_{adv}^{T_{adv}}=KL(I(x_{adv}),T_{adv}(x_{adv})),$ (7)
where $I$ denotes the general model with random weights.
Our goal is to reduce the gap between clean teachers’ knowledge scale $K_{nat}^{T_{nat}}$ and robust teachers’ knowledge scale $K_{adv}^{T_{adv}}$, and we define the gap as $\Delta_{K}$. To further simplify the gap $\Delta_{K}$, a theorem 1 is provided, and the theorem 1 is proved in proof 1.
Theorem 1
The knowledge scale’s difference between clean and robust teacher $\Delta_{K}$ can be expressed by the difference between the teacher’s information entropy $H(P_{nat}^{T_{nat}})$ and $H(P_{adv}^{T_{adv}})$:
$\Delta_{K}=||H(P_{adv}^{T_{adv}})-H(P_{nat}^{T_{nat}})||.$ (8)
Proof 1
Based on information theory, the difference in teachers’ relative entropy in knowledge distillation is as follows:
$\Delta_{K}=||K_{adv}^{T_{adv}}-K_{nat}^{T_{nat}}||.$ (9)
For a model with randomly initialized weights, the model’s prediction has no preference for any samples, so a reasonable assumption can be made: at an initial state, when the general model $I$ outputs $P_{nat}^{I}=\{p_{1}^{I}(x_{nat}),...,p_{C}^{I}(x_{nat})\}$ for $x_{nat}$ and $P_{adv}^{I}=\{p_{1}^{I}(x_{adv}),...,p_{C}^{I}(x_{adv})\}$ for $x_{adv}$, thus:
$p_{k}^{I}(x_{nat})=p_{k}^{I}(x_{adv})=1/C,k=1,2,...,C.$ (10)
Assume the clean teacher’s predicted distribution for clean examples is $P_{nat}^{T_{nat}}$, and the robust teacher’s predicted distribution for adversarial examples is $P_{adv}^{T_{adv}}$, we further expand the $\Delta_{K}$ as follows:
$\Delta_{K}$ $=||H(P_{adv}^{T_{adv}})-\sum_{k=1}^{C}p_{k}^{T_{adv}}(x_{adv})log(p_{k}^{I}(x_{adv}))$
$+\sum_{k=1}^{C}p_{k}^{T_{nat}}(x_{nat})log(p_{k}^{I}(x_{nat}))-H(P_{nat}^{T_{nat}})||,$ (11)
based on the assumption of the initial state (Eq. (10)), we have:
$\Delta_{K}=$ $||H(P_{adv}^{T_{adv}})-log(1/C)\sum_{k=1}^{C}p_{k}^{T_{adv}}(x_{adv})$
$+log(1/C)\sum_{k=1}^{C}p_{k}^{T_{nat}}(x_{nat})-H(P_{nat}^{T_{nat}})||,$ (12)
$=$ $||H(P_{adv}^{T_{adv}})-H(P_{nat}^{T_{nat}})||.$ (13)
Then the Theorem 1 is proved.
Based on Theorem 1, by reducing the difference in the teachers’ information entropy, teachers’ knowledge scales will tend to be consistent, which is conducive to mitigating the trade-off between accuracy and robustness. So we define the above knowledge scales gap as an optimization objective in mathematical form as follows:
$\min\Delta_{K}=||H(P_{adv}^{T_{adv}})-H(P_{nat}^{T_{nat}})||.$ (14)
To solve the optimization object $\min\Delta_{K}$, the intuitive strategy is to jointly re-train the teachers so that the information entropy is relatively consistent. However, the training process requires an additional calculation consumption, which is not an ideal approach.
Then the hyper-parameter temperature $\tau$ attracts our attention. To properly utilize the teacher, a common practice in the field of knowledge distillation is adjusting the temperature $\tau$, which can directly influence the teacher’s predicted distribution and produce a suitably soft label for the student. Two characteristics of the temperature make us believe that adjusting the temperature can achieve the optimization goal elegantly. Firstly, the adjustment of the temperature can directly transform the information entropy of the model’s prediction, and the effect is as shown in Fig. 2: When the temperature changes, the prediction’s information entropy also changes obviously. Secondly, the calculation cost of updating temperature is almost negligible, which means that we may achieve the final goal if a good optimization towards temperature $\tau$ is designed.
Figure 2: Schematic diagram of the same logits but dealing with different temperatures $\tau$ and getting different predicted probabilities $p$. From the figure, the temperature $\tau$ can impact the model prediction and can further impact the information entropy $H$.
Based on the above consideration, we expect to dynamically update the multiple teachers’ temperatures in the training process to achieve our final optimization goal. More specially, our optimization goal has changed from balancing the teacher’s knowledge scales to indirectly adjusting the information entropy by adjusting the temperature.
To solve the above optimization Eq. (14), we need to apply the gradient of the expected object $\Delta_{K}$ concerning the temperature $\tau_{nat}$ and $\tau_{adv}$ and update them according to the gradient descent principle in the training process. Here we apply the optimization process of temperature $\tau_{nat}$ to expand our method (the optimization process of temperature $\tau_{adv}$ is similar, and we do not repeat the description), which is formulated as follows:
$\tau_{nat}=\tau_{nat}-r_{\tau}\nabla_{\tau_{nat}}\Delta_{K},$ (15)
we use the learning rate $r_{\tau}$ to control the magnitude of the gradient change, and then we further expand the preliminary derivation:
$\begin{split}\nabla_{\tau_{nat}}\Delta_{K}=\left\{\begin{array}[]{ll}\nabla_{\tau_{nat}}H(P_{nat}^{T_{nat}}),&H(P_{nat}^{T_{nat}})\geq H(P_{adv}^{T_{adv}}),\\ -\nabla_{\tau_{nat}}H(P_{nat}^{T_{nat}}),&H(P_{nat}^{T_{nat}})<H(P_{adv}^{T_{adv}}).\end{array}\right.\end{split}$ (16)
Here, we calculate the partial derivative of information entropy $H(P)$ ($P$ can be considered as a general expression including $P_{nat}^{T_{nat}}$ and $P_{adv}^{T_{adv}}$, and $\tau$ is the same) with respect to temperature $\tau$ and can get the result as follows:
$\nabla_{\tau}H(P)=\frac{(\sum_{j=1}^{C}q_{j})(\sum_{j=1}^{C}q_{j}log^{2}q_{j})-(\sum_{j=1}^{C}q_{j}logq_{j})^{2}}{\tau(\sum_{j=1}^{C}q_{j})^{2}},$ (17)
where the $q$ represents the $exp(z(x)/\tau)$ in Eq. (3), and the more detailed derivation can be found in Appendix A (Proof A.1.). Meanwhile, we prove that the partial derivative $\nabla_{\tau}H(P)\geq 0$, and this conclusion is proved in Appendix A (Proof A.2.), thus the information entropy $H(P)$ shows a monotonically increasing relationship with the temperature $\tau$. Based on the consideration of the stable update, we use the constant 1 to estimate $\nabla_{\tau}H(P)$ in the training process of updating the teacher’s temperature, and we use $r_{\tau}$ to control the update range.
Then combined the above proof with Eq. (15) and Eq. (16), the formula for updating the $\tau_{nat}$ is as follows:
$\tau_{nat}=\tau_{nat}-r_{\tau}sign(H(P_{nat}^{T_{nat}})-H(P_{adv}^{T_{adv}})),$ (18)
where $sign(.)$ denotes the sign function. According to the same derivation process, we can also obtain the formula for updating the temperature $\tau_{adv}$ for teacher $T_{adv}$ can be formulated as follows:
$\tau_{adv}=\tau_{adv}-r_{\tau}sign(H(P_{adv}^{T_{adv}})-H(P_{nat}^{T_{nat}})).$ (19)
On a practical level, our Entropy-Based Balance algorithm can balance the teacher’s knowledge scale by controlling the information entropy. When the information entropy of one teacher is greater or smaller than another teacher, both teachers’ temperatures will be updated until the information entropy of the two teachers reaches a relatively balanced state, and the knowledge scales of different teachers will be equally matched. With the adjustment of the Entropy-Based Balance, the clean teacher and the robust teacher can appear equally knowledgeable when imparting knowledge to the student, which is helpful to mitigate the trade-off between accuracy and robustness.
It should be mentioned that we only update the teachers’ temperatures ($\tau_{nat}$ and $\tau_{adv}$) and keep the student’s temperature ($\tau_{s}$) unchanged in the Entropy-Based Balance. In addition, in order to prevent information from being over-regulated, we set an upper and lower bound for the temperature to prevent the following situations. If the temperature is too small, the teacher’s predicted distribution will be infinitely close to the one-hot label’s distribution; When the temperature is too large, the teacher’s predicted distribution will be close to the uniform distribution.
3.3 Normalization Loss Balance in MTARD
Although the two teachers are equally knowledgeable after the adjustment of the Entropy-Based Balance algorithm, the student may not learn equal knowledge from the two teachers during the training process, since the difficulties of recognizing clean examples and adversarial examples for the student are different: Adversarial examples are generated to mislead the student based on clean examples. In order to let the student get equal knowledge from the clean teacher and the robust teacher, a strategy is needed to control the student’s relative learning speed toward different types of knowledge. Inspired by gradient regularization methods in multi-task learning [4], we propose an algorithm to control the relative speed by dynamically adjusting the loss weights in the entire training process, which is called the Normalization Loss Balance algorithm.
On the mathematical level, the total loss in MTARD ultimately used for the student update at time $t$ can be represented as $L_{total}(t)$, which can be formulated as follows:
$L_{total}(t)=w_{nat}(t)L_{nat}(t)+w_{adv}(t)L_{adv}(t),$ (20)
where the $L_{nat}(t)$ and $L_{adv}(t)$ denote the KL loss mentioned in Eq. (4), and the $w_{nat}(t)$ and $w_{adv}(t)$ denote the constant hyper-parameters $1-\alpha$ and $\alpha$ mentioned in Eq. (4) in our task if without the Normalization Loss Balance. Similar to the analysis in section 3.2, the knowledge scales’ gap between the student and the teacher can be directly reflected by the value of the clean loss $L_{nat}(t)$ and adversarial loss $L_{adv}(t)$, and the vital to control the student’s knowledge from multiple teachers consistent is controlling the loss weight of $w_{nat}(t)$ and $w_{adv}(t)$, which directly influence the student’s learning of $L_{nat}(t)$ and $L_{adv}(t)$.
Then our goal is to place $L_{nat}(t)$ and $L_{adv}(t)$ on a common metric through their relative magnitudes. Then based on the common metric, we can identify specific optimization directions to ensure that $L_{nat}(t)$ and $L_{adv}(t)$ have a relatively fair descent after the entire update process, and the final trained model can be equally affected by different parts of the total loss.
To select the criterion metric to measure the decline of multiple losses, we choose a relative loss $\tilde{L}(t)$ following Chen et al. [4], which is defined as follows:
$\tilde{L}_{nat}(t)=L_{nat}(t)/L_{nat}(0),~{}\tilde{L}_{adv}(t)=L_{adv}(t)/L_{adv}(0),$ (21)
especially in our task, $L_{nat}(t)$ and $L_{adv}(t)$ represent the knowledge scales’ gap for the student from the clean teacher $T_{nat}$ and robust teacher $T_{adv}$ at time $t$. So $\tilde{L}_{nat}(t)$ and $\tilde{L}_{adv}(t)$ can reflect the knowledge proportion that the student has not learned from the different teacher’s knowledge at the time $t$ compared with time 00. When $\tilde{L}_{nat}(t)$ and $\tilde{L}_{adv}(t)$ are equal, the student obtains relatively equal knowledge scales from different teachers, which is exactly what we search for. So our optimization goal is as follows:
$\min||\tilde{L}_{nat}(t)-\tilde{L}_{adv}(t)||.$ (22)
During the student’s training process, the most direct impact on $\tilde{L}_{nat}(t)$ and $\tilde{L}_{adv}(t)$ is the loss weight $w_{nat}$ and $w_{adv}$. Intuitively speaking, when $\tilde{L}_{nat}(t)$ is smaller than $\tilde{L}_{adv}(t)$, the student has learned more knowledge scales from the clean teacher compared with the robust teacher during the 0 to t time, and the student is supposed to have smaller loss weight $w_{nat}$ in the learning process to limit the learning speed from this type of knowledge.
Based on the above criterion and analysis, we can dynamically balance $\tilde{L}_{nat}(t)$ and $\tilde{L}_{adv}(t)$ by applying a relative weight $r_{nat}(t)$ and $r_{adv}(t)$ that we expected for loss at time t, which can be formulated as follows:
$r_{nat}(t)=[\tilde{L}_{nat}(t)]^{\beta}/\big{[}[\tilde{L}_{nat}(t)]^{\beta}+[\tilde{L}_{adv}(t)]^{\beta}\big{]},$ (23)
$r_{adv}(t)=[\tilde{L}_{adv}(t)]^{\beta}/\big{[}[\tilde{L}_{nat}(t)]^{\beta}+[\tilde{L}_{adv}(t)]^{\beta}\big{]},$ (24)
where $[\tilde{L}(t)]^{\beta}$ denotes the $\tilde{L}(t)$ power of $\beta$, and $\beta$ is set to control the strength of correcting the imbalance. The $r_{nat}(t)$ and $r_{adv}(t)$ strengthen the disadvantaged losses and weakens the advantaged losses to eliminate the imbalance between the student’s learning speeds in different types of knowledge. The big $\beta$ is applicable when the vibration of relative loss value is obvious, while the small $\beta$ is suitable for $\tilde{L}_{nat}(t)$ and $\tilde{L}_{adv}(t)$ with stable influence abilities.
From the consideration of training stability, we do not directly use $r_{nat}(t)$ and $r_{adv}(t)$ as the weight of the loss, but update the original weight $w_{nat}(t)$ and $w_{adv}(t)$, which can be formulated as follows:
$w_{nat}(t)=r_{w}r_{nat}(t)+(1-r_{w})w_{nat}(t-1),$ (25)
$w_{adv}(t)=r_{w}r_{adv}(t)+(1-r_{w})w_{adv}(t-1).$ (26)
On a practical level, with the assistance of the Normalization Loss Balance algorithm, the student can dynamically adjust the different loss weights in the whole training process according to the degree of acceptance of knowledge. The student’s learning speed will be inhibited and the corresponding loss weight will decrease if too much of this type of knowledge is accepted; On the contrary, the student’s learning speed will be accelerated and the corresponding loss weight will increase if this type of knowledge is insufficient. Finally, the student tends to learn well from both two teachers, gaining relatively equal clean and robust knowledge, rather than appearing to have a partial ability.
3.4 Overview of MTARD
All in all, to transfer both clean and robust knowledge to the student, we propose the MTARD to apply two teachers and propose the Entropy-Based Balance algorithm for teachers and the Normalization Loss Balance algorithm for the student to optimize the adversarial training process. The final minimization of MTARD is defined as follows:
$\begin{split}\mathop{\arg\min}_{\theta_{S}}&w_{nat}KL(S(x_{nat};\tau_{s}),T_{nat}(x_{nat};\tau_{nat}))\\ &+w_{adv}KL(S(x_{adv};\tau_{s}),T_{adv}(x_{adv};\tau_{adv})),\end{split}$ (27)
and the complete process of MTARD is in Algorithm 1.
Algorithm 1 Overview of MTARD
1:Initialize student $S$ with weight $\theta_{S}$ and temperature $\tau_{s}$, pretrained teacher $T_{nat}$ and $T_{adv}$, the initial temperature $\tau_{nat}$ and $\tau_{adv}$ for $T_{nat}$ and $T_{adv}$, respectively, the benign clean examples $x_{nat}$ and the labels $y$, the threat bound $\Omega$, the max training epochs $max$-$epoch$.
2:for $epoch=0$ to $max$-$epoch$ do
3:     for $t=0$ to $max$-$step$ do
4:         Acquire adversarial example $x_{adv}$ by Eq. (5).
5:         Get $L_{nat}(t)=KL(S(x_{nat};\tau_{s}),T_{nat}(x_{nat};\tau_{nat}))$.
6:         Get $L_{adv}(t)=KL(S(x_{adv};\tau_{s}),T_{adv}(x_{adv};\tau_{adv}))$.
7:         if $epoch=0$ and $t=0$ then
8:              Record $L_{nat}$ and $L_{adv}$ as $L_{nat}(0)$ and $L_{adv}(0)$.
9:         end if
10:         Update $w_{nat}(t)$ and $w_{adv}(t)$ by Eq. (25) and Eq. (26).
11:         Update $\theta_{S}$ by Eq. (27).
12:         Update $\tau_{nat}$ and $\tau_{adv}$ by Eq. (18) and Eq. (19).
13:     end for
14:end for
Compared with other existing methods, our method has several advantages. Firstly, our method can fit the different teacher-student combinations, which can be automatically adjusted by the Entropy-Based Balance algorithm and Normalization Loss Balance algorithm. Secondly, our method pays more attention to the overall performance based on Weighted Robust Accuracy, which measures the trade-off between accuracy and robustness. Thirdly, the training hyper-parameters can be dynamically updated by our adjustment algorithm and can fit the changes as training epochs increase, which is important to fit various scenarios.
4 Experiments
In this section, we initially describe the experimental setting and then evaluate the Weighted Robust Accuracy of our MTARD and several baseline methods under prevailing white-box attacks and black-box attacks including transfer-based and square-based attacks. We also conduct a series of ablation studies to demonstrate the effectiveness of our method. Our code is available at https://github.com/zhaoshiji123/MTARD-extension.
4.1 Experimental Settings
We conduct our experiments on three datasets including CIFAR-10 [17], CIFAR-100, and Tiny-ImageNet [19], and consider the standard training method and several state-of-the-art methods as comparison methods: Adversarial Training method: SAT [25], Adversarial Robustness Distillation methods: ARD [9] and RSLAD [49]; Mitigating Trade-off methods: TRADES [45] and SCORE [26].
TABLE I: Performance of the teacher networks used in our experiments.
Dataset
Model
Clean Acc
FGSM
[…]
${}_{\scriptscriptstyle\rm{\infty}}$
AA
Type
CIFAR-10
ResNet-56
93.18%
19.23%
0%
0%
0%
0%
Clean
WideResNet-34-10
84.91%
61.14% 55.30% 56.61% 53.84% 52.53%
Adv
CIFAR-100
WideResNet-22-6
76.65%
4.85%
0%
0%
0%
0%
Clean
WideResNet-70-16
60.96%
35.89% 33.58% 33.99% 31.05% 30.03%
Adv
Tiny-ImageNet
PreActResNet-34
66.05%
2.50%
0.04%
0.09%
0%
0%
Clean
PreActResNet-34
52.76%
27.05% 24.00% 24.63% 20.07% 18.92%
Adv
Student and Teacher Networks. Here we consider two student networks for CIFAR-10 and CIFAR-100 including ResNet-18 [13] and MobileNet-v2 [28]. For Tiny-ImageNet, we select PreActResNet-18 [14] and MobileNet-v2 following previous work. As for teacher models, we choose clean teacher networks including ResNet-56 for CIFAR-10, WideResNet-22-6 [44] for CIFAR-100, and PreActResNet-34 [14] for Tiny-ImageNet. Three robust teacher networks include WideResNet-34-10 for CIFAR-10, WideResNet-70-16 [11] for CIFAR-100, and PreActResNet-34 for Tiny-ImageNet. For CIFAR-10, WideResNet-34-10 is trained using TRADES [45]; For CIFAR-100, we use the WideResNet-70-16 model provided by Gowal et al. [11]; In Addition, for Tiny-ImageNet, we use PreActResNet-34 trained by TRADES [45] as the robust teacher. The whole teachers are pre-trained and will not be changed during the training process. The performance of these teacher models is shown in Table I.
TABLE II: White-box robustness on CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets. The results of CIFAR-10 and CIFAR-100 are based on the ResNet-18, and the results of Tiny-ImageNet are based on the PreActResNet-18. All results are the best checkpoint based on W-Robust Acc.
CIFAR-10 CIFAR-100 Tiny-ImageNet
Attack
Defense
Clean
Robust
W-Robust
Clean
Robust
W-Robust
Clean
Robust
W-Robust
FGSM[10]
Natural
94.57%
18.60%
56.59%
75.18%
7.96%
41.57%
65.03%
1.73%
33.38%
SAT[25]
84.20%
55.59%
69.90%
56.16%
25.88%
41.02%
50.08%
25.35%
37.72%
TRADES[45]
83.00%
58.35%
70.68%
57.75%
31.36%
44.56%
48.45%
23.59%
36.02%
ARD[9]
84.11%
58.40%
71.26%
60.11%
33.61%
46.86%
53.22%
27.97%
40.60%
RSLAD[49]
83.99%
60.41%
72.20%
58.25%
34.73%
46.49%
48.78%
27.26%
38.02%
SCORE[26]
84.43%
59.84%
72.14%
56.40%
32.94%
44.67%
10.05%
7.80%
8.93%
MTARD(ours)
88.20%
61.42%
74.81%
65.08%
34.21%
49.65%
56.81%
28.12%
42.47%
[…]
Natural
94.57%
0%
47.29%
75.18%
0%
37.59%
65.03%
0.03%
32.53%
SAT[25]
84.20%
45.95%
65.08%
56.16%
21.18%
38.67%
50.08%
22.24%
36.16%
TRADES[45]
83.00%
52.35%
67.68%
57.75%
28.05%
42.90%
48.45%
21.59%
35.02%
ARD[9]
84.11%
50.93%
67.52%
60.11%
29.4%%
44.76%
53.22%
24.92%
39.07%
RSLAD[49]
83.99%
53.94%
68.97%
58.25%
31.19%
44.72%
48.78%
25.00%
36.89%
SCORE[26]
84.43%
53.72%
69.08%
56.40%
30.27%
43.34%
10.05%
7.65%
8.85%
MTARD(ours)
88.20%
51.68%
69.94%
65.08%
28.50%
46.79%
56.81%
23.93%
40.37%
[…]
Natural
94.57%
0%
47.29%
75.18%
0%
37.59%
65.03%
0.05%
32.54%
SAT[25]
84.20%
48.12%
66.16%
56.16%
22.02%
39.09%
50.08%
23.05%
36.57%
TRADES[45]
83.00%
53.83%
68.42%
57.75%
28.88%
43.32%
48.45%
22.09%
35.27%
ARD[9]
84.11%
52.96%
68.54%
60.11%
30.51%
45.31%
53.22%
25.71%
39.47%
RSLAD[49]
83.99%
55.73%
69.86%
58.25%
32.05%
45.15%
48.78%
25.45%
37.12%
SCORE[26]
84.43%
55.21%
69.82%
56.40%
30.56%
43.48%
10.05%
7.67%
8.86%
MTARD(ours)
88.20%
54.40%
71.30%
65.08%
29.94%
47.51%
56.81%
24.94%
40.88%
${}_{\scriptscriptstyle\rm{\infty}}$[3]
Natural
94.57%
0%
47.29%
75.18%
0%
37.59%
65.03%
0%
32.52%
SAT[25]
84.2%
45.97%
65.09%
56.16%
20.90%
38.53%
50.08%
20.48%
35.28%
TRADES[45]
83.00%
50.23%
66.62%
57.75%
24.19%
40.97%
48.45%
17.33%
32.89%
ARD[9]
84.11%
50.15%
67.13%
60.11%
27.56%
43.84%
53.22%
21.41%
37.32%
RSLAD[49]
83.99%
52.67%
68.33%
58.25%
28.21%
43.23%
48.78%
20.87%
34.83%
SCORE[26]
84.43%
50.46%
67.45%
56.40%
26.30%
41.35%
10.05%
6.19%
8.13%
MTARD(ours)
88.20%
49.88%
69.04%
65.08%
25.45%
45.27%
56.81%
19.69%
38.25%
AutoAttack[6]
Natural
94.57%
0%
47.29%
75.18%
0%
37.59%
65.03%
0%
32.52%
SAT[25]
84.20%
43.49%
63.85%
56.16%
19.76%
37.96%
50.08%
18.54%
34.31%
TRADES[45]
83.00%
49.03%
66.02%
57.75%
23.26%
40.51%
48.45%
16.66%
32.56%
ARD[9]
84.11%
48.20%
66.16%
60.11%
25.71%
42.91%
53.22%
19.91%
36.57%
RSLAD[49]
83.99%
50.98%
67.49%
58.25%
26.50%
42.38%
48.78%
17.11%
32.95%
SCORE[26]
84.43%
49.25%
66.84%
56.40%
24.74%
40.57%
10.05%
5.97%
8.01%
MTARD(ours)
88.20%
47.44%
67.82%
65.08%
23.98%
44.53%
56.81%
17.71%
37.26%
TABLE III: White-box robustness on CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets. The result of three datasets is trained based on the MobileNet-v2. All results are the best checkpoint based on W-Robust Acc.
CIFAR-10 CIFAR-100 Tiny-ImageNet
Attack
Defense
Clean
Robust
W-Robust
Clean
Robust
W-Robust
Clean
Robust
W-Robust
FGSM[10]
Natural
93.35%
12.22%
52.79%
74.86%
5.94%
40.40%
62.09%
0.72%
31.41%
SAT[25]
83.87%
55.89%
69.88%
59.19%
30.88%
45.04%
49.03%
23.38%
36.21%
TRADES[45]
77.95%
53.75%
65.85%
55.41%
30.28%
42.85%
43.81%
20.10%
31.96%
ARD[9]
83.43%
57.03%
70.23%
60.45%
32.77%
46.61%
45.53%
22.88%
33.21%
RSLAD[49]
83.20%
59.47%
71.34%
59.01%
33.88%
46.45%
45.69%
24.09%
34.89%
SCORE[26]
82.32%
58.43%
70.38%
49.38%
29.28%
39.33%
28.27%
17.47%
22.87%
MTARD(ours)
88.70%
58.78%
73.74%
66.13%
34.36%
50.25%
52.98%
25.60%
39.29%
[…]
Natural
93.35%
0%
46.68%
74.86%
0%
37.43%
62.09%
0%
31.05%
SAT[25]
83.87%
46.84%
65.36%
59.19%
25.64%
42.42%
49.03%
20.31%
34.67%
TRADES[45]
77.95%
49.06%
63.51%
55.41%
23.33%
39.37%
43.81%
18.16%
30.99%
ARD[9]
83.43%
49.50%
66.47%
60.45%
28.69%
44.57%
45.53%
20.43%
32.98%
RSLAD[49]
83.20%
53.25%
68.23%
59.01%
30.19%
44.60%
45.69%
22.30%
34.00%
SCORE[26]
82.32%
53.42%
67.87%
49.38%
27.03%
38.21%
28.27%
16.48%
22.38%
MTARD(ours)
88.70%
48.50%
68.60%
66.13%
28.47%
47.30%
52.98%
21.58%
37.28%
[…]
Natural
93.35%
0%
46.68%
74.86%
0%
37.43%
62.09%
0%
31.05%
SAT[25]
83.87%
49.14%
66.51%
59.19%
26.96%
43.08%
49.03%
21.15%
35.09%
TRADES[45]
77.95%
50.27%
64.11%
55.41%
28.42%
41.92%
43.81%
18.36%
31.09%
ARD[9]
83.43%
51.70%
67.57%
60.45%
29.63%
45.04%
45.53%
21.00%
33.27%
RSLAD[49]
83.20%
54.76%
68.98%
59.01%
31.19%
45.10%
45.69%
22.74%
34.22%
SCORE[26]
82.32%
54.46%
68.39%
49.38%
27.53%
38.46%
28.27%
16.69%
22.48%
MTARD(ours)
88.70%
51.11%
69.91%
66.13%
29.82%
47.98%
52.98%
22.58%
37.78%
${}_{\scriptscriptstyle\rm{\infty}}$[3]
Natural
93.35%
0%
46.68%
74.86%
0%
37.43%
62.09%
0%
31.05%
SAT[25]
83.87%
46.62%
65.25%
59.19%
25.01%
42.10%
49.03%
18.69%
33.86%
TRADES[45]
77.95%
46.06%
62.01%
55.41%
27.72%
41.57%
43.81%
13.47%
28.66%
ARD[9]
83.43%
48.96%
66.20%
60.45%
26.55%
43.50%
45.53%
16.81%
31.17%
RSLAD[49]
83.20%
51.78%
67.49%
59.01%
27.98%
43.50%
45.69%
18.63%
32.16%
SCORE[26]
82.32%
49.18%
65.75%
49.38%
23.29%
36.34%
28.27%
13.25%
20.76%
MTARD(ours)
88.70%
47.09%
67.90%
66.13%
26.50%
46.32%
52.98%
18.08%
35.53%
AutoAttack[6]
Natural
93.35%
0%
46.68%
74.86%
0%
37.43%
62.09%
0%
31.05%
SAT[25]
83.87%
43.43%
63.65%
59.19%
22.84%
41.02%
49.03%
17.35%
33.19%
TRADES[45]
77.95%
45.18%
61.66%
55.41%
22.61%
39.01%
43.81%
12.82%
28.32%
ARD[9]
83.43%
46.60%
65.02%
60.45%
24.86%
42.66%
45.53%
15.40%
30.47%
RSLAD[49]
83.20%
50.23%
66.72%
59.01%
26.33%
42.67%
45.69%
17.09%
31.39%
SCORE[26]
82.32%
48.39%
65.36%
49.38%
21.92%
35.65%
28.27%
12.30%
20.29%
MTARD(ours)
88.70%
44.93%
66.82%
66.13%
24.01%
45.07%
52.98%
15.68%
34.33%
Training Setting. We train the student using Stochastic Gradient Descent (SGD) optimizer with an initial learning rate of 0.1, a momentum of 0.9, and a weight decay of 2e-4. For MTARD, the weight loss learning rate $r_{w}$ is initially set as 0.025; meanwhile, the $\tau_{nat}$ and $\tau_{adv}$ are initially set as 1, and the temperature learning rate $r_{\tau}$ is initially set as 0.001. the student’s temperature $\tau_{s}$ is set to constant 1. The maximum and minimum values of temperature are 10 and 1 respectively. For CIFAR-10 and CIFAR-100, we set the total number of training epochs to 300. The learning rate is divided by 10 at the 215-th, 260-th, and 285-th epochs; For Tiny-ImageNet, we set the total number of training epochs to 100, and the learning rate is divided by 10 at the 75-th and 90-th epochs. We set the batch size to 128 (CIFAR-10 and CIFAR-100) and 32 (Tiny-ImageNet), and $\beta$ is set to 1. For the inner maximization of MTARD, we use a 10-step PGD with a random start size of 0.001 and a step size of 2/255. All training perturbation in the maximization is bounded to the $L_{\infty}$ norm $\epsilon$ = 8/255. For the training of the natural model, we train the networks for 100 epochs, and the learning rate is divided by 10 at the 75-th and 90-th epochs.
For the comparison methods, we follow SAT, TRADES, SCORE, ARD, and RSLAD original settings if without additional instruction. The SCORE version is the TRADES+LSE without additional data. For ARD, we use the same robust teacher as RSLAD and MTARD. The temperature $\tau$ of ARD is set to 30 on CIFAR-10 while set to 5 on CIFAR-100 and Tiny-ImageNet, and the $\alpha$ is set to 0.95 following [9] on CIFAR-100 and Tiny-ImageNet. The learning rate of SAT and SCORE for Tiny-ImageNet is 0.01 to achieve better performance.
[…]
Figure 3: Ablation study with ResNet-18 student trained using variants of our MTARD and Baseline method on CIFAR-10. NLB and EBB represent Normalization Loss Balance (NLB) and Entropy-Based Balance (EBB). MTARD is our final version, which represents Baseline+NLB+EBB. All the results are the best checkpoint based on W-Robust Acc.
Figure 4: The total loss $L_{total}$ with the ResNet-18 student trained using variants of RSLAD, Baseline, and Baseline+NLB on CIFAR-10. NLB is abbreviation of Normalization Loss Balance. The y-axis is the $L_{total}$ in the training epoch x. The left is the change curve of $L_{total}$ in the whole training process, the right is the curve of $L_{total}$ in the final 60 epochs.
Here, we use Weighted Robust Accuracy (W-Robust Acc) [12] to evaluate the trade-off between the clean and robust accuracy of the model, it is defined as follows:
$\mathcal{A}_{f}=\pi_{{nat}}P_{{nat}}[f(x)=y]+\pi_{{adv}}P_{{adv}}[f(x)=y],$ (28)
where W-Robust Acc $\mathcal{A}_{f}$ are the accuracy of a model $f$ on x drawn from either the clean distribution $P_{nat}$ and the adversarial distribution $P_{adv}$. We set $\pi_{{nat}}$ and $\pi_{{adv}}$ both to 0.5, which means accuracy and robustness are equally important for comprehensive performance in the model.
4.2 Ablation Study
Figure 5: The relative training loss $\tilde{L}$ with ResNet-18 student on CIFAR-10. The left is Baseline, and the right is Baseline+NLB. NLB represents Normalization Loss Balance. The x-axis means the training epochs, and the y-axis is the relative loss $\tilde{L}_{adv}$ and $\tilde{L}_{nat}$ in the training epoch x.
Figure 6: The teachers’ information entropy of $T_{adv}$ and $T_{nat}$ with ResNet-18 student on CIFAR-10. The left is Baseline+NLB, and the right is MTARD (Baseline+NLB+EBB). NLB and EBB are abbreviations of Normalization Loss Balance and Entropy-based Balance. The x-axis means the training epochs (the first 40 epochs), and the y-axis is the entropy $T_{adv}$ and $T_{nat}$ in the training epoch x.
Figure 7: The result with different initial temperature settings in MTARD. The left is the results of different ranges for teachers’ temperatures. The right is the results of different initial values for teachers’ temperatures. The initial values of $\tau_{nat}$ and $\tau_{adv}$ for $T_{adv}$ and $T_{nat}$ are equal. All the results are the best checkpoint based on W-Robust Acc.
TABLE IV: Study with different $\beta$, the students are ResNet-18 trained on CIFAR-10 guided by ResNet-56 and WideResNet-34-10. All the results are the best checkpoint based on W-Robust Acc.
Attack
$\beta$ value
Clean
Robust
W-Robust
[…]
$\beta=0.5$
88.51%
50.24%
69.38%
$\beta=1$
88.20%
51.68%
69.94%
$\beta=4$
89.10%
50.01%
69.56%
$\beta=10$
88.63%
49.09%
68.86%
[…]
$\beta=0.5$
88.51%
53.43%
70.97%
$\beta=1$
88.20%
54.40%
71.30%
$\beta=4$
89.10%
53.04%
71.07%
$\beta=10$
88.63%
52.63%
70.63%
${}_{\scriptscriptstyle\rm{\infty}}$[3]
$\beta=0.5$
88.51%
48.59%
68.55%
$\beta=1$
88.20%
49.88%
69.04%
$\beta=4$
89.10%
48.61%
68.86%
$\beta=10$
88.63%
48.16%
68.40%
To better understand the impact of each component in our MTARD, we conduct a set of ablation studies. The baseline denotes using a clean teacher and a robust teacher to guide the student from clean and robust aspects, respectively, where the weight $w_{nat}$ and $w_{adv}$ are constant at 0.5. Baseline+NLB term denotes adding the Normalization Loss Balance (NLB) algorithm to dynamically adjust the weight $w_{nat}$ and $w_{adv}$ based on the Baseline. Baseline+EBB term denotes adding the Entropy-Based Balance (EBB) algorithm to dynamically adjust the $\tau_{nat}$ and $\tau_{adv}$ based on the Baseline. MTARD is the final version of our method. The performance is shown in Fig. 3. The change of total loss $L_{total}$ is shown in Fig. 4, and the change of relative loss $\tilde{L}_{nat}$ and $\tilde{L}_{adv}$ is shown in Fig. 5. Meanwhile, the change for the teacher’s information entropy is shown in Fig. 6.
In Fig. 3, MTARD’s improvement brought by each component is remarkable, which shows its effectiveness. Multiple teachers positively affect the student model to learn with both clean and robust accuracy. However, it is not enough to mitigate the trade-off between accuracy and robustness without the Entropy-based Balance algorithm and the Normalization Loss Balance algorithm.
TABLE V: Black-box robustness on CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets. The results of CIFAR-10 and CIFAR-100 are based on the ResNet-18, and the results of Tiny-ImageNet are based on the PreActResNet-18. All results are the best checkpoint based on W-Robust Acc.
CIFAR-10 CIFAR-100 Tiny-ImageNet
Attack
Defense
Clean
Robust
W-Robust
Clean
Robust
W-Robust
Clean
Robust
W-Robust
[…]
SAT[25]
84.20%
64.84%
74.52%
56.16%
38.10%
47.13%
50.08%
33.40%
41.74%
TRADES[45]
83.00%
63.61%
73.31%
57.75%
38.20%
47.98%
48.45%
31.01%
39.73%
ARD[9]
84.11%
63.50%
73.81%
60.11%
39.53%
49.82%
53.22%
34.74%
43.98%
RSLAD[49]
83.99%
63.96%
73.98%
58.25%
39.93%
49.09%
48.78%
32.85%
40.82%
SCORE[26]
84.43%
64.34%
74.39%
56.40%
39.09%
47.75%
10.05%
8.74%
9.40%
MTARD(ours)
88.20%
65.29%
76.75%
65.08%
42.11%
53.60%
56.81%
36.65%
46.73%
${}_{\scriptscriptstyle\rm{\infty}}$[3]
SAT [25]
84.20%
63.84%
74.02%
56.16%
39.42%
47.79%
50.08%
33.20%
41.63%
TRADES[45]
83.00%
62.83%
72.92%
57.75%
38.63%
48.19%
48.45%
30.72%
39.59%
ARD[9]
84.11%
62.86%
73.49%
60.11%
38.85%
49.48%
53.22%
33.32%
43.27%
RSLAD[49]
83.99%
63.05%
73.52%
58.25%
39.67%
48.96%
48.78%
32.09%
40.44%
SCORE[26]
84.43%
63.79%
74.11%
56.40%
40.02%
48.21%
10.05%
8.82%
9.44%
MTARD(ours)
88.20%
64.64%
76.42%
65.08%
41.35%
53.22%
56.81%
33.40%
45.11%
Square Attack[1]
SAT [25]
84.20%
72.48%
78.34%
56.16%
40.05%
48.11%
50.08%
38.72%
44.40%
TRADES[45]
83.00%
72.49%
77.75%
57.75%
42.51%
50.13%
48.45%
36.58%
42.52%
ARD[9]
84.11%
74.60%
79.36%
60.11%
47.20%
53.66%
53.22%
42.58%
47.90%
RSLAD[49]
83.99%
72.47%
78.23%
58.25%
45.32%
51.79%
48.78%
37.64%
43.21%
SCORE[26]
84.43%
74.68%
79.56%
56.40%
43.98%
50.34%
10.05%
8.67%
9.36%
MTARD(ours)
88.20%
79.82%
84.01%
65.08%
49.40%
57.24%
56.81%
44.46%
50.64%
TABLE VI: Black-box robustness on CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets. The result of three datasets is trained based on the MobileNet-v2. All results are the best checkpoint based on W-Robust Acc.
CIFAR-10 CIFAR-100 Tiny-ImageNet
Attack
Defense
Clean
Robust
W-Robust
Clean
Robust
W-Robust
Clean
Robust
W-Robust
[…]
SAT [25]
83.87%
64.66%
74.27%
59.19%
40.70%
49.95%
49.03%
33.47%
41.25%
TRADES[45]
77.95%
61.04%
69.50%
55.41%
37.76%
46.59%
43.81%
28.35%
36.08%
ARD[9]
83.43%
63.28%
73.36%
60.45%
39.15%
49.80%
45.53%
30.73%
38.13%
RSLAD[49]
83.20%
64.33%
73.77%
59.01%
40.32%
49.67%
45.69%
31.20%
38.45%
SCORE[26]
82.32%
63.84%
73.08%
49.38%
40.00%
44.69%
28.27%
21.82%
25.05%
MTARD(ours)
88.70%
66.20%
77.45%
66.13%
42.67%
54.40%
52.98%
34.25%
43.62%
${}_{\scriptscriptstyle\rm{\infty}}$[3]
SAT [25]
83.87%
64.24%
74.06%
59.19%
40.97%
50.08%
49.03%
33.13%
41.08%
TRADES[45]
77.95%
60.66%
69.31%
55.41%
38.02%
46.72%
43.81%
28.64%
36.23%
ARD[9]
83.43%
62.83%
73.13%
60.45%
38.53%
49.49%
45.53%
30.23%
37.88%
RSLAD[49]
83.20%
63.45%
73.33%
59.01%
39.92%
49.47%
45.69%
31.10%
38.40%
SCORE[26]
82.32%
63.03%
72.68%
49.38%
36.91%
43.15%
28.27%
22.19%
25.23%
MTARD(ours)
88.70%
65.73%
77.22%
66.13%
42.04%
54.09%
52.98%
33.50%
43.24%
Square Attack[1]
SAT [25]
83.87%
73.01%
78.44%
59.19%
44.51%
51.85%
49.03%
37.95%
43.49%
TRADES[45]
77.95%
67.43%
72.69%
55.41%
40.71%
48.06%
43.81%
32.39%
38.10%
ARD[9]
83.43%
73.26%
78.35%
60.45%
46.95%
53.70%
45.53%
34.60%
40.07%
RSLAD[49]
83.20%
73.07%
78.14%
59.01%
45.66%
52.34%
45.69%
35.18%
40.44%
SCORE[26]
82.32%
72.41%
77.37%
49.38%
37.94%
43.66%
28.27%
22.16%
25.22%
MTARD(ours)
88.70%
79.30%
84.00%
66.13%
50.83%
58.48%
52.98%
40.62%
46.80%
In Fig. 4, compared with RSLAD and Baseline, the Baseline+NLB’s training loss is less oscillating and can converge better. In Fig. 5, the gap between $\tilde{L}_{nat}$ and $\tilde{L}_{adv}$ can represent the difference of the knowledge proportion that students have not learned from the clean teacher and the robust teacher. In the final training period, the Baseline+NLB’s trade-off between $\tilde{L}_{nat}$ and $\tilde{L}_{adv}$ are tinier. The results demonstrate Normalization Loss Balance can indeed make the student have relatively equal knowledge learning speeds from different teachers.
In Fig. 6, we can see that in Baseline+NLB, the information entropy gap between the clean teacher and the robust teacher is very obvious, but after the adjustment of our Entropy-based Balance algorithm, the information entropy gap quickly narrows and remains relatively consistent, which proves the effectiveness of our Entropy-based Balance algorithm. Combined with the experimental results in Fig. 3, reducing the difference in teacher’s information entropy can indeed improve the performance of the student.
4.3 The discussion about hyper-parameters
Hyper-parameter $\tau$. The temperatures of different teachers are automatically adjusted by our Entropy-based algorithm until the information entropy is equal. However, the initial temperature setting still directly influences the amount of information entropy for both teachers, which has an inevitable impact on the final performance. Intuitively speaking, the temperature is supposed to change in a proper range to maintain the effectiveness of teachers’ knowledge (If the temperature is close to infinity, the predicted distribution is close to uniform distribution; If the temperature is close to 0, the predicted distribution is close to one-hot distribution). In order to explore the proper range for temperature, we select three ranges: [0.1, 1], [1, 10], and [10, 20] for teachers’ temperatures to change (The initial temperatures for different teachers are median of the ranges: 0.5, 5, 15, respectively). Based on the result shown in Fig. 7 (left), we select the [1, 10] as our initial setting. Meanwhile, to select the proper initial temperature value, we select initial $\tau_{nat}$ and $\tau_{adv}$ as 1, 2, 5, 8 in the range [1, 10]. Based on the result shown in Fig. 7 (right), we select 1 as our initial temperature value.
Hyper-parameter $\beta$. Here, we explore the role of $\beta$ in Normalization Loss Balance (Eq. (23) and Eq. (24)). We choose ResNet-56 and WideResNet-34-10 as the clean and robust teachers, respectively, and select ResNet-18 as the student on CIFAR-10. We test the results of different values $\beta$ for the training of the student. We select $\beta$ as 0.5, 1, 4, and 10 in our experiment. The result is presented in Table IV.
From the result, $\beta$ plays an important role in Normalization Loss Balance. Choosing the right $\beta$ value can affect the student’s final performance. The lower $\beta$ is suitable when the student gets a similar knowledge scale from the clean teacher and the robust teacher at the current training epoch and the student’s status is not fluctuating much. The larger $\beta$ brings a more severe penalty on the student when the student places too much emphasis on one teacher’s knowledge but ignores another. Based on the performance, we choose the relatively best value: $\beta=1$ in our setting.
4.4 Adversarial Robustness Evaluation
White-box Robustness. The performances of ResNet-18 and MobileNet-v2 trained by our MTARD and other baseline methods under the white box attacks are shown in Table II and III for CIFAR-10, CIFAR-100, and Tiny-ImageNet. We select the best checkpoint of the baseline model and MTARD based on Weighted Robust Accuracy.
[…]
Black-box Robustness. In addition, we also test MTARD and other methods against black-box attacks for ResNet-18 and MobileNet-v2 on CIFAR-10, CIFAR-100, and Tiny-ImageNet separately. We choose the transfer-based attack and query-based attack in our evaluation. We select the best checkpoint of all the trained models based on Weighted Robust Accuracy. The result of ResNet-18 is shown in Table V, while the result of MobileNet-v2 is shown in Table VI.
[…]
TABLE VII: The gradient estimation (GE) Attack [33] toward the ResNet-18 trained by MTARD on different datasets. The results are the robust Acc.
Attack
CIFAR-10
CIFAR-100
Tiny-ImageNet
[…]
54.40%
29.94%
24.94%
GE attack [33]
60.34%
32.53%
30.83%
Excluding Obfuscated Gradients. Although some defense methods claim to resist adversarial attacks, their vulnerabilities are often exposed because they belong to “Obfuscated Gradients” [33, 2]. Athalye et al. [2] have confirmed that adversarial training [25] can effectively defend against the adaptive attack and do not belong to “Obfuscated Gradients”. Here we argue that MTARD is excluding the “Obfuscated Gradients” from the several aspects following [2] and [35]: (1) Our MTARD can effectively defend against strong attack: AutoAttack [6], which includes strong white-box attack: A-PGD and strong black-box attack: Square Attack [1]. (2) Strong white-box test attack (e.g. ${}_{\scriptscriptstyle\rm{\infty}}$[3]) have higher attack success rates than weak white-box test attack (e.g. FGSM [10]) in Table II and Table III. (3) White-box test attacks have higher attack success rates than Black-box test attacks (Comparing with Table II and Table V, Table III and Table VI). (4) To avoid insufficient attack rounds and to fall into the local optimal solution in the general attack methods, a gradient estimation attack method from [33] is directly applied to test MTARD. The result in Table VII shows MTARD can effectively resist the strong gradient estimation attack. Meanwhile, the black-box attack success rate is lower than white-box attack success rate, which further confirms MTARD is not “Obfuscated Gradients”.
4.5 Further Explorations for Different Teacher Models
To better understand the impact of different teachers, we evaluate the performance of the same students guided by different teachers. We choose ResNet-18 as the student, and ResNet-56 and WideResNet-34-10-C (95.83% clean accuracy) as the clean teacher, respectively. Also, we choose WideResNet-34-10 and WideResNet-70-16 (From [11] with stronger robustness) as the robust teacher, and other experimental settings are the same as the initial settings. The performance is shown in Table VIII.
TABLE VIII: Study with different teacher models using our MTARD, WRN-34-10-C is the clean teacher model WideResNet-34-10-C, WRN-34-10 and WRN-70-16 are the abbreviations of WideResNet-34-10 and WideResNet-70-16. RN-56 is the abbreviation of ResNet-56. All the results are the best checkpoints based on W-Robust Acc.
Attack
Clean Teacher
Robust Teacher
W-Robust
[…]
RN-56
WRN-34-10
69.94%
WRN-34-10-C
WRN-34-10
69.09%
WRN-34-10-C
WRN-70-16
68.49%
${}_{\scriptscriptstyle\rm{\infty}}$[3]
RN-56
WRN-34-10
69.04%
WRN-34-10-C
WRN-34-10
68.48%
WRN-34-10-C
WRN-70-16
67.89%
To our surprise, the student’s ability is not improved as the teacher’s parameter scale increases. Although larger teachers have stronger accuracy and robustness, the student guided by clean ResNet-56 and adversarial WideResNet-34-10 achieves the best W-Robust Acc instead of other students guided by larger teachers, which demonstrates that small teachers still have pretty potential to guide the student and can bring effective improvement to the student. Meanwhile, the performance of the student is not strictly bound to the performance of the teacher and the model size, which still has potential space to explore.
5 Conclusion
This paper focused on mitigating the trade-off between accuracy and robustness in adversarial training. To bring both clean and robust knowledge, we proposed Multi-Teacher Adversarial Robustness Distillation (MTARD) to guide the student model, in which a clean teacher and a robust teacher are applied during the adversarial training process. In addition, we proposed the Entropy-Based Balance algorithm to keep the knowledge scale of these teachers consistent. In order to ensure that the student gets equal knowledge from two teachers, we designed a method to use Normalization Loss Balance in MTARD. A series of experiments demonstrated that our MTARD outperformed the existing adversarial training methods and adversarial robustness distillation on CIFAR-10, CIFAR-100, and Tiny-ImageNet. In the future, MTARD can be applied to other tasks with multiple optimization goals, not only limited to adversarial robustness distillation, which has great development potential.